On a number line, point A has a coordinate of -6, and point B has a coordinate of 2. Which is the coordinate of point M, the midpoint of AB?
A)4
B)-3
C)0
D)-2

Answer by richwmiller(17219)

6+2=8
8/2=4
-6+4=-2
d)-2